# Ground-Compose-Reinforce: Tasking Reinforcement Learning Agents through Formal Language

### v1

@@ 4 Problem Setting
We aim to develop RL agents that are fully taskable through formal language: given only a formal language specification, it must learn to solve the corresponding task by interacting with the environment. We avoid manually designing reward functions or labelling functions that capture task-relevant aspects of the environment—while these methods conveniently help ground language in the environment, they become increasingly hard to specify as tasks and environments grow in complexity. However, without a way to ground language, any language-based specification is meaningless to an agent acting in an environment. Imagine, for example, that your boss tells you to “get the dax, then throw away the blick[66, 67] but you don’t know what “dax” and “blick” mean.
We explore how formal language can be grounded by training on a fixed grounding dataset, and then subsequently applied for specifying RL tasks. We formulate this problem in the context of RMs.
Problem Formulation. Let $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mu,\gamma\rangle$ be an MDP without rewards or terminal states, and let $\mathcal{AP}$ be a finite set of propositions. We assume a ground-truth labelling function $\mathcal{L}^{*}:\mathcal{S}\to 2^{\mathcal{AP}}$ exists, but is unobservable. Instead, we assume we have a fixed grounding dataset $\mathcal{D}=\{\langle\tau^{i},\omega^{i}\rangle\}_{i=1}^{N}$ of trajectories $\tau^{i}=\langle s^{i}_{0},a^{i}_{0},s^{i}_{1},\ldots\rangle$ with corresponding labels $\omega^{i}=\langle\mathcal{L}^{*}(s^{i}_{0}),\mathcal{L}^{*}(s^{i}_{1}),\ldots\rangle$. We aim to use $\mathcal{D}$ to develop an agent that can learn to solve any RM task $\mathcal{R}=\langle\mathcal{U},u_{0},\mathcal{F},\mathcal{AP},\delta_{u},\delta_{r}\rangle$ over $\mathcal{AP}$ without supervision. That is, it can learn a policy $\pi_{\mathcal{R}}(a_{t}|h_{t})$ that performs well in the corresponding RM-MDP $\langle\mathcal{M},\mathcal{R},\mathcal{L}^{*}\rangle$ by interacting with $\mathcal{M}$, but without access to RM-MDP rewards or $\mathcal{L}^{*}$. [^1] For the problem statement, we consider the more general form of history-based policies, rather than policies conditioned on the ground-truth RM state, which depends on $\mathcal{L}^{*}$.
Running Example.
Suppose our gridworld agent is expected to solve any new RM task over the vocabulary $\{\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}},\mathrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}G}},\mathrm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}},\triangle,\bigcirc\}$. The agent may interact freely with the environment, but observes neither ground-truth rewards nor propositional labels (e.g. whether $\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}}$ holds in the current state) while doing so. Instead, it is first given a fixed set $\mathcal{D}$ of trajectories labelled with this vocabulary—from this, it must learn how to interpret any such RM task in the context of the environment.
Since the agent faces arbitrary RM tasks without reward supervision, it must acquire a grounded understanding of both individual symbols (e.g., ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}$) and their compositions as RM tasks to reliably solve new tasks. We also consider sparse settings in which $\mathcal{L}^{*}(s)$ is constant over most of the reachable state space under random exploration. This is common in domains like robotics, where random exploration rarely produces meaningfully diverse or goal-relevant behaviour. We hypothesize that existing RM-based approaches fail to learn effectively under such conditions.
@@ 5 Ground-Compose-Reinforce
We address the problem through a framework called Ground-Compose-Reinforce. At a high level, the agent first grounds the symbols of a formal language by training on $\mathcal{D}$. Given any task expressed in this language, it can then evaluate trajectories by leveraging both its grounded interpretation of symbols and the compositional semantics of the formal language (i.e., how symbol meanings combine to form complex tasks). Finally, the agent learns to solve new tasks through RL without supervision, instead using its ability to self-evaluate behaviours to generate its own learning signals.
We hypothesize that grounding language bottom-up—first learning individual symbols, then composing them to interpret complex tasks—is more effective than learning to associate entire task specifications with behaviours in an end-to-end fashion. We test this hypothesis in our experiments.
We now describe the core implementation of Ground-Compose-Reinforce for RMs. This version supports fully unsupervised RL for new RM tasks. In Section 6, we extend this framework to handle sparse settings by learning a compact set of value functions from $\mathcal{D}$ and using them to approximate value functions for any RM task. We show that this enables the agent to solve long-horizon tasks without sacrificing optimality on the target task.
@@
0:  MDP $\mathcal{M}$ without rewards, Propositional symbols $\mathcal{AP}$, Dataset $\mathcal{D}$ of labelled trajectories, RM task $\mathcal{R}$ over $\mathcal{AP}$ // Phase 1: Ground symbols via offline learning
1:  Train labelling function ^$\hat{\mathcal{L}}(s)$ on $\mathcal{D}$ using any binary classifier // Phase 2: Solve any RM task using self-generated rewards
@@
9:        Update RM: $u^{\prime}\leftarrow\delta_{u}(u,\hat{\omega})$, $r\leftarrow\delta_{r}(u,\hat{\omega})$
@@
12:     end while
13:  end for
Core Algorithm. We first ground the propositional symbols $\mathcal{AP}$ by training a model of the labelling function ^$\hat{\mathcal{L}}(s)\approx\mathcal{L}^{*}(s)$ on $\mathcal{D}$ using any binary classification method. Then, given any RM task $\mathcal{R}$ over $\mathcal{AP}$, we solve a surrogate RM-MDP $\langle\mathcal{M},\mathcal{R},\hat{\mathcal{L}}\rangle$ via RL. This surrogate task approximates the true RM-MDP $\langle\mathcal{M},\mathcal{R},\mathcal{L}^{*}\rangle$, for which the agent lacks supervision. Although the agent cannot access $\mathcal{L}^{*}$ or ground-truth rewards, it can query $\hat{\mathcal{L}}$ freely—thus, it can simulate rewards $\langle{r}_{1},{r}_{2},\ldots\rangle$ and RM states $\langle{u}_{1},{u}_{2},\ldots\rangle$ for any trajectory, as described in Section 3.3. We use these self-generated signals to train a policy $\pi_{\mathcal{R}}(a_{t}|s_{t},u_{t})$ for the equivalent MDP over state space $\mathcal{S}\times\mathcal{U}$. We outline this full process in Algorithm 1.
@@
With Ground-Compose-Reinforce, the gridworld agent first learns to interpret propositions $\{\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}},\mathrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}G}},\mathrm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}},\triangle,\bigcirc\}$ in relation to its environment, grounding each proposition to states via the dataset $\mathcal{D}$. Once grounded, the agent can understand any RM task (e.g. those in Figure 1) that composes these symbols through temporal or logical structure. A human can then specify a new RL task as an RM $\mathcal{R}$, without needing to program a task-specific reward function. The agent learns by predicting its own rewards and RM transitions based on its interpretation of the task.
@@ 6 Compositional Reward Shaping
We now address the challenge of sparsity by extending Ground-Compose-Reinforce with potential-based reward shaping [68], using approximations of optimal value functions (OVFs) as shaping potentials. Specifically, we estimate the OVF for the surrogate RM-MDP $\langle\mathcal{M},\mathcal{R},\hat{\mathcal{L}}\rangle$ by decomposing $\mathcal{R}$ into simpler subtasks. The most basic subtasks correspond to satisfying individual propositions or their negations, and we estimate the OVFs for these $2|\mathcal{AP}|$ tasks directly from $\mathcal{D}$ using offline RL. These $2|\mathcal{AP}|$ OVFs then serve as building blocks that can be composed to approximate the OVF for any of the infinitely many RM tasks over $\mathcal{AP}$. We describe this compositional shaping method below.
@@ 6.1 Deconstructing RMs into Logical Subtasks
To estimate the OVF of the surrogate RM-MDP $\langle\mathcal{M},\mathcal{R},\hat{\mathcal{L}}\rangle$, denoted $V^{*}_{\mathcal{R}}(s,u)$, we combine two key ideas: (1) treating each RM transition in $\mathcal{R}$ as an independent subtask, and (2) bootstrapping from a state-independent value function $v^{*}_{\mathcal{R}}(u)$.
Consider the example in Figure 2, where the agent is in MDP state $s$ and RM state $u^{A}$. Each RM transition is associated with a propositional condition $\varphi$ (e.g., ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}\land\triangle$ or ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}\land\lnot\triangle$), and we treat the satisfaction of $\varphi$ as its own subtask. Definition 3 formalizes this: we introduce the class $\Diamond\mathbf{PL}(\mathcal{AP})$ of reachability tasks over propositional formulas, where each task $\Diamond\varphi$ entails reaching a state $s_{T}$ such that ^$\hat{\mathcal{L}}(s_{T})\models\varphi$, terminating with reward 1 upon satisfaction. $\Diamond\varphi$ is itself an RM-MDP with a single non-terminal RM state, and is therefore Markovian over $\mathcal{S}$. We denote its OVF as $V^{*}_{\Diamond\varphi}(s)$.
@@
For any propositional logic formula $\varphi$ over $\mathcal{AP}$, define $\Diamond\varphi$ as the task of reaching a state $s_{T}\in\mathcal{S}$ such that ^$\hat{\mathcal{L}}(s_{T})\models\varphi$. The episode terminates and yields a reward of 1 if such a state is reached, and continues indefinitely with 0 reward otherwise. Let $\Diamond\mathbf{PL}(\mathcal{AP})=\allowbreak\left\{\Diamond\varphi\;\middle|\;\varphi\text{ is a propositional formula over }\mathcal{AP}\right\}$ denote the set of such reachability tasks.
@@
This approximation assumes no RM self-transitions with non-zero rewards. Appendix A provides further justification for our approximation and also extends it to handle any RM.
@@ 6.2 Further Deconstructing Logical Subtasks
Approximation 1 allows us to estimate $V^{*}_{\mathcal{R}}(s,u)$ for any RM task $\mathcal{R}$ over $\mathcal{AP}$, provided we can estimate $V^{*}_{\Diamond\varphi}(s)$ for any $\varphi$ in $\Diamond\mathbf{PL}(\mathcal{AP})$. However, the number of propositional formulas over $\mathcal{AP}$ (up to logical equivalence) is $2^{2^{|\mathcal{AP}|}}$ and modelling a separate OVF for each such task is intractable. To address this, we further decompose logical formulas based on their structure. Any formula $\varphi$ can be rewritten in disjunctive normal form, i.e., as a disjunction of conjunctions of literals (where a literal is either a proposition $x$ or its negation $\lnot x$). We then approximate the OVF of $\Diamond\varphi$ using the semantics of fuzzy logic [69], where $\max$ represents disjunction and $\min$ represents conjunction. [^2] While fuzzy operators have been widely applied in formal languages over quantitative signals [51, 52, 53, 54], we apply them here to approximate OVFs. To the best of our knowledge, this has not been previously explored.
@@ 6.3 Final Remarks
In this section, we showed that the optimal value function (OVF) of any RM task $\mathcal{R}$ can be approximated using just $2|\mathcal{AP}|$ primitive value functions (PVFs). From these $2|\mathcal{AP}|$ PVFs, we can estimate OVFs for doubly exponentially many logical tasks ($2^{2^{n}}$) and infinitely many RM tasks. Each PVF quantifies progress toward satisfying a single proposition or its negation, and can be learned directly from the grounding dataset $\mathcal{D}$ using any offline RL algorithm. This approach trades expressivity for modularity: directly modelling the OVFs of all RM tasks is infeasible, so we instead model a small, reusable set of $2|\mathcal{AP}|$ PVFs at the cost of introducing approximation error.
Leveraging that we can estimate the OVF for any RM task $\mathcal{R}$, we then extend Algorithm 1 via potential-based reward shaping. This allows us to provide a dense learning signal, even in environments that suffer from sparsity. Further details on learning PVFs, sources of approximation error, and the potential-based reward shaping scheme are provided in Appendix A.
Running Example.
Suppose our gridworld agent learns OVFs for the propositions $\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}},\mathrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}G}},\mathrm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}},\triangle,\bigcirc$ and their negations. For propositions, these OVFs quantify progress toward reaching, for example, a red object or a circle. For negations, the OVFs capture progress toward reaching a state where the given proposition is false—e.g., any square without a red object. Note that these negations can typically be satisfied in a single step and their OVFs are trivial to model. These $2|\mathcal{AP}|$ OVFs are the primitive value functions (PVFs)—from these, the agent can estimate the OVF for any RM task.
@@ 1. RQ1
    Grounded (Formal) Language Learning: Can formal language be grounded more effectively by leveraging compositional language structure?
@@ 2. RQ2
    Compositional Generalization: Do grounded language interpretations generalize to out-of-distribution tasks beyond those observed in $\mathcal{D}$?

### v2

@@ 4 Problem Setting
Our goal in this work is to faithfully elicit behaviours from an agent given only a high-level task specification in the form of an RM, $\mathcal{R}$ such as the ones depicted in Figure 2. $\mathcal{R}$ can be specified directly, translated from other formal languages like LTL [34], generated from a symbolic planner [66, 67], or sometimes autoformalized from natural language. Formally, we consider an environment $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mu,\gamma\rangle$ (an MDP without rewards or terminal states) and a finite set of propositional symbols $\mathcal{AP}$. For any given RM task $\mathcal{R}=\langle\mathcal{U},u_{0},\mathcal{F},\mathcal{AP},\delta_{u},\delta_{r}\rangle$, we wish to obtain a policy $\pi_{\mathcal{R}}(a_{t}|h_{t})$ that performs well in the RM-MDP $\langle\mathcal{M},\mathcal{R},\mathcal{L}^{*}\rangle$, where $\mathcal{L}^{*}:\mathcal{S}\to 2^{\mathcal{AP}}$ reflects a ground-truth interpretation of the propositions $\mathcal{AP}$ in the environment. [^2] For the problem statement, we consider the more general form of history-based policies, rather than policies conditioned on the ground-truth RM state, which depend on $\mathcal{L}^{*}$.
Assumptions. We aim to obtain $\pi_{\mathcal{R}}$ without online access to $\mathcal{L^{*}}$ or to an external reward function that evaluates ground-truth performance with respect to $\mathcal{R}$. [^3] Such oracles can be notoriously hard to design in practice [68] and often require internal simulator access. In order to connect symbols $\mathcal{AP}$ with environment percepts, we instead assume access to a fixed pretraining dataset $\mathcal{D}=\{\langle\tau^{i},\omega^{i}\rangle\}_{i=1}^{N}$ of trajectories $\tau^{i}=\langle s^{i}_{0},a^{i}_{0},s^{i}_{1},\ldots\rangle$ with corresponding labels $\omega^{i}=\langle\mathcal{L}^{*}(s^{i}_{0}),\mathcal{L}^{*}(s^{i}_{1}),\ldots\rangle$. In practice, such labels can be obtained via crowdsourced annotations [69] or self-supervised learning [70].
For a task $\mathcal{R}$, the agent is allowed an arbitrary number of interaction episodes with the environment before committing to a final policy $\pi_{\mathcal{R}}$. However, during this interaction phase, the agent must learn in a self-supervised manner as the environment does not provide a separate reward signal.
@@ 5 Ground-Compose-Reinforce
We propose an end-to-end framework for this setting called Ground-Compose-Reinforce (Figure 1). In the pretraining phase, the agent first grounds propositional symbols $\mathcal{AP}$ in environment states via supervised learning on $\mathcal{D}$. In the behaviour elicitation phase, the agent is given a task as an RM $\mathcal{R}$ composed over symbols $\mathcal{AP}$. The agent then learns a policy $\pi_{\mathcal{R}}$ by interacting with the environment and synthesizing its own learning signal for RL based on $\mathcal{R}$ and its learned interpretation of $\mathcal{AP}$. We hypothesize that this bottom-up approach to grounding language in behaviours—first learning the meanings of individual symbols and then composing them to interpret complex tasks—is an effective strategy. The remainder of this section describes the core implementation of Ground-Compose-Reinforce and in Section 6, we raise and address an issue called propositional sparsity where the agent fails to learn in extremely long-horizon tasks.
@@
0: MDP $\mathcal{M}$ without rewards, Propositional symbols $\mathcal{AP}$, Dataset $\mathcal{D}$ of labelled trajectories, RM task $\mathcal{R}$ over $\mathcal{AP}$ // Pretraining phase
1: Train labelling function $\hat{\mathcal{L}}(s)$ on $\mathcal{D}$ using any binary classification method // Behaviour elicitation phase
@@
[…]
@@
Core Algorithm. We ground the propositional symbols $\mathcal{AP}$ by learning a labelling function $\hat{\mathcal{L}}(s)\approx\mathcal{L}^{*}(s)$ via any binary classification method on $\mathcal{D}$. Given an RM task $\mathcal{R}$ over $\mathcal{AP}$, we solve a surrogate RM-MDP $\langle\mathcal{M},\mathcal{R},\hat{\mathcal{L}}\rangle$ via RL. This surrogate task approximates the true RM-MDP $\langle\mathcal{M},\mathcal{R},\mathcal{L}^{*}\rangle$ for which the agent lacks supervision. Since the agent can query $\hat{\mathcal{L}}$ freely, it can simulate rewards $\langle{r}_{1},{r}_{2},\ldots\rangle$ and RM states $\langle{u}_{1},{u}_{2},\ldots\rangle$ for any trajectory, as described in Section 3.3. Finally, we use these self-generated signals to train a policy $\pi_{\mathcal{R}}(a_{t}|s_{t},u_{t})$ as outlined in Algorithm 1.
@@
[…]
@@ 6 Compositional Reward Shaping
An issue that can arise with the core implementation is sparse rewards. While prior RM works have proposed additional learning signals for transitioning in the RM [34, 48, 49], these methods fail when RM transitions themselves pose an exploration challenge. This can be caused by propositional sparsity: when $\mathcal{L}^{*}(s)$ is constant across most states under random exploration, the agent struggles to learn to meaningfully affect propositional values (e.g. consider a robot manipulation task that involves satisfying the proposition “the box is picked up”).
To address this, we extend Ground-Compose-Reinforce with potential-based reward shaping [71]. Specifically, for a given RM task $\mathcal{R}$, we estimate the optimal value function (OVF) for the surrogate RM-MDP $\langle\mathcal{M},\mathcal{R},\hat{\mathcal{L}}\rangle$ by decomposing $\mathcal{R}$ into simpler subtasks. The most basic subtasks correspond to satisfying individual propositions or their negations, and we estimate the OVFs for these $2|\mathcal{AP}|$ tasks during pretraining via offline RL on $\mathcal{D}$. These $2|\mathcal{AP}|$ OVFs then serve as building blocks that can be composed to approximate the OVF for any of the infinitely many RM tasks over $\mathcal{AP}$.
@@ 6.1 Deconstructing RMs into Logical Subtasks
To estimate the OVF of the surrogate RM-MDP $\langle\mathcal{M},\mathcal{R},\hat{\mathcal{L}}\rangle$, denoted $V^{*}_{\mathcal{R}}(s,u)$, we combine two key ideas: (1) treating each RM transition in $\mathcal{R}$ as an independent subtask, and (2) bootstrapping from a state-independent value function $v^{*}_{\mathcal{R}}(u)$. Consider the example in Figure 3, where the agent is in MDP state $s$ and RM state $u^{A}$. Each RM transition is associated with a logical condition $\varphi$ (e.g., $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R$$\land\triangle$ or $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B$$\land\lnot\triangle$), and we treat the satisfaction of $\varphi$ as its own subtask. Definition 3 formalizes this: we introduce the class $\Diamond\mathbf{PL}(\mathcal{AP})$ of reachability tasks over propositional formulas, where each task $\Diamond\varphi$ entails reaching a state $s_{T}$ such that $\hat{\mathcal{L}}(s_{T})\models\varphi$, terminating with reward 1 upon satisfaction. $\Diamond\varphi$ is itself an RM-MDP with a single non-terminal RM state (and a terminal RM state), and is therefore Markovian over $\mathcal{S}$. We denote its OVF as $V^{*}_{\Diamond\varphi}(s)$.
@@
For any propositional logic formula $\varphi$ over $\mathcal{AP}$, define $\Diamond\varphi$ as the task of reaching a state $s_{T}\in\mathcal{S}$ such that $\hat{\mathcal{L}}(s_{T})\models\varphi$. The episode terminates and yields a reward of 1 if such a state is reached, and continues indefinitely with 0 reward otherwise. Let $\Diamond\mathbf{PL}(\mathcal{AP})=\allowbreak\left\{\Diamond\varphi\;\middle|\;\varphi\text{ is a propositional formula over }\mathcal{AP}\right\}$ denote the set of such reachability tasks.
@@
This approximation assumes no RM self-transitions with non-zero rewards. Appendix A further justifies and explains our approximation while extending it to arbitrary RMs.
@@ 6.2 Further Deconstructing Logical Subtasks
Approximation 1 allows us to estimate $V^{*}_{\mathcal{R}}(s,u)$ for any RM task $\mathcal{R}$ over $\mathcal{AP}$, provided we can estimate $V^{*}_{\Diamond\varphi}(s)$ for any $\varphi$ in $\Diamond\mathbf{PL}(\mathcal{AP})$. However, the number of propositional formulas over $\mathcal{AP}$ (up to logical equivalence) is $2^{2^{|\mathcal{AP}|}}$ and modelling a separate OVF for each such task is intractable. To address this, we further decompose logical formulas based on their structure. Any formula $\varphi$ can be rewritten in disjunctive normal form, i.e., as a disjunction of conjunctions of literals (where a literal is either a proposition $x$ or its negation $\lnot x$). We then approximate the OVF of $\Diamond\varphi$ using the semantics of fuzzy logic [72], where $\max$ represents disjunction and $\min$ represents conjunction. [^4] Fuzzy operators have previously been applied for satisfaction of a temporal formula over quantitative signals [33, 53, 54]. While such tasks are binary in nature, we consider RMs, which can express other reward structures.
@@ 6.3 Final Remarks
In this section, we showed that the optimal value function (OVF) of any RM task $\mathcal{R}$ can be approximated using just $2|\mathcal{AP}|$ primitive value functions (PVFs). From these $2|\mathcal{AP}|$ PVFs, we can estimate OVFs for doubly exponentially many logical tasks ($2^{2^{n}}$) and infinitely many RM tasks. Each PVF quantifies progress toward satisfying a single proposition or its negation, and can be learned directly from $\mathcal{D}$ using any offline RL algorithm. One might view this approach as trading off expressivity for modularity: directly modelling the OVFs of all RM tasks is infeasible, so we instead model a small, reusable set of $2|\mathcal{AP}|$ PVFs at the cost of introducing some approximation error.
Leveraging that we can estimate the OVF for any RM task $\mathcal{R}$, we extend Algorithm 1 with potential-based reward shaping to address propositional sparsity. Further details on learning PVFs, sources of approximation error, and the potential-based reward shaping scheme are provided in Appendix A.
@@ 1. RQ1
    Grounding RMs in Behaviours: With Ground-Compose-Reinforce, can we faithfully elicit behaviours given high-level task specifications (RMs)?
@@ 2. RQ2
    Compositional Generalization: Can we elicit meaningful, out-of-distribution (OOD) behaviours beyond those observed in $\mathcal{D}$?